OpenStudy (anonymous):

@ganeshie8 @uri Cards numbered 1, 1, 2, 3, 4, 4, 4, 5, 6, 6, 7, 8, 8, 8, and 9 are placed in a bag. If a card is drawn from the bag and then replaced 135 times, about how many times would a 6 be drawn? A. 2 B. 9 C. 15 D. 18

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

do u know how to do this

OpenStudy (texaschic101):

well.....there are 15 cards...and 2 of them are 6....so the chance of drawing a 6 on one try is 2/15.....so I guess on 135 tries, it is 2/15 * 135 = 270/15 = 18 times. I am not the best at this...do you think this looks right ?

OpenStudy (anonymous):

ya it looks right i so the answer is D.

OpenStudy (texaschic101):

I think so :)

OpenStudy (anonymous):

do u know anyone u can tag

OpenStudy (texaschic101):

hold on...let me see who is on

OpenStudy (texaschic101):

Its not letting me tag....type in @jim_thompson or @phi or @jdoe0001. They will know

OpenStudy (texaschic101):

oh..it worked on 2 of them

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (texaschic101):

I think it is right..logically, it looks right

OpenStudy (texaschic101):

I got to go now....sorry...I will be back on later.

OpenStudy (anonymous):

kk bye thanks